Very low seismic activity continued during the period under review. Several small long-period earthquakes (magnitude less than 1), were recorded by the nearest seismometer stations (Chances Peak and Gages). These events were too small to be located and varied in their rate of occurrence throughout the day. Approximately 20 other seismic signals, of the type that can often be associated with rockfall or small explosions were also large enough to be observed on the four nearest stations to the volcano. Many more of these signals are only observed on the new station at Chances Peak but are to small to be seen even on the previous closest seismometer at Gages.
A new station was installed at Chances Peak on 8 February and after some adjustment yesterday is now fully operational. This station is now the most sensitive to activity in English's Crater and along with the recently established station at Hermitage will improve the quality of earthquake locations.
An attempt to measure the Dagenham-Gages-Amersham triangle failed today due to low cloud cover and mist in the Gages valley. No measurements were made on the eastern triangle to Castle Peak and low cloud also prevented measurement of the Chances Peak triangle.
Visual observation of the volcano was made from the helicopter on two occasions today and from Farrells early this afternoon. Brief views of activity on the western parts of the dome confirmed observations made from Chances Peak yesterday. The focus of dome growth has shifted to the western parts of the new dome.. Several rockfalls were seen and heard to occur on the western parts of the dome. Brief views were obtained of a prominent spine which has been extruded in the western part of the dome during the past week. Observations from the crater wall at Farrells reveal that the northern part of the dome has stopped growing. Vigorous steaming was observed from the top central parts of the dome and from the south eastern parts of the active area.
The light blue haze which has been visible in the Gages valley for the past week shifted today and no longer extend beyond the upper parts of this valley.
Mr Desmond Seupersad returned to the Seismic Research Unit in Trinidad yesterday after spending three weeks on the island assisting with processing of seismic signals and with EDM measurements. Mr Seupersad was replaced by Mr Lutchman Pollard who arrived this morning from Trinidad. Mr Pollard would assist in the processing and analysis of seismic events.